# Efficient and accurate force replay in cosmological-baryonic simulations

### v1

@@ Abstract
journal: APJsoftware: IPython (Perez & Granger, 2007), Matplotlib (Hunter, 2007), Numpy (Harris et al., 2020), Gizmo Analysis (Wetzel & Garrison-Kimmel, 2020a), Agama (Vasiliev, 2019b), Rockstar (Behroozi et al., 2012), Halo Analysis (Wetzel & Garrison-Kimmel, 2020b), CMasher (van der Velden, 2020).
@@ 1 Introduction
The time-dependent evolution of the halo structure can be captured by using Basis Function Expansions (BFE) models, which can accurately describe any arbitrary density field (e.g Weinberg, 1999; Lowing et al., 2011; Vasiliev, 2013; Sanders et al., 2020; Garavito-Camargo et al., 2021; Vasiliev et al., 2021; Arora et al., 2022; Petersen et al., 2022; Lilleengen et al., 2023). Lowing et al. (2011) showed that by applying a Hernquist BFE (often referred to in the literature as a self-consistent field expansion) to every snapshot, spaced roughly 25 Myr, from a cosmological simulation of a DM halo and then interpolating the coefficients of the expansion in time allows substantial improvement in the orbital properties of a halo. Similarly, Sanders et al. (2020) showed high-fidelity orbit reconstruction in DM-only cosmological simulations of MW mass galaxies using compact BFE representation for the time-evolving halo potential, employing spherical harmonics for angular variation and bi-orthonormal basis functions or splines for radial variation.
In this paper, we extend this methodology to model fully zoomed baryonic-cosmological simulations of MW mass galaxies from the Latte suite of the FIRE-2 project (Wetzel et al., 2016, 2023). Unlike Sanders et al. (2020), where only DM was considered, our simulations include stars and gas, necessitating adaptation of the BFE method with an azimuthal harmonic expansion to account for the baryonic component’s flattened shape. Additionally, we utilize a fiducial temporal cadence with snapshots spaced at approximately 25 Myr intervals, and explore the effect of larger sampling intervals (up to 500 Myr) on the quality of recovered orbits. We demonstrate that sufficiently high fidelity halo orbit reconstruction can be achieved with our fiducial cadence, even though the potential in the inner regions can change much more rapidly. We assess the efficacy of this modified potential model in reconstructing orbits at a range of radii. We fit BFE to the potential at a series of discrete snapshots. Then, we approximate the time dependence by calculating a force/acceleration on a star at a given point in time as a linear interpolation of the forces in the two time-adjacent snapshots for orbit recovery.
@@
In Sec. 2, we provide a brief overview of the simulation (§2.1), potential modeling techniques (§2.2), constraints on recovered particle forces in the potential model (§2.3), and describe our sample of selected halo stars for orbit integration (§2.4). In Sec. 3, we statistically quantify the quality of reconstructed orbits based on recovered 3D positions (§3.1), approximate integrals of motion, and orbital parameters such as pericenter and apocenter distances (§3.2). We also evaluate the dependence of orbit quality on the sampling interval for potential models (§3.4). In Sec. 4, we demonstrate an application of the model by simulating the tidal disruption of a dwarf satellite. We discuss our findings and conclusions in Sec. 5.
@@ 2 Methods
In this section, we detail our approach to reconstructing halo star orbits in a simulation of a MW-mass galaxy. We describe the galaxy simulation used (Sec. 2.1), the BFE-based potential model fits to the simulation (Sec. 2.2) and their force reconstruction for a sample of DM and star particles from the parent halo at present day (Sec. 2.3). We also describe how we select stellar orbits from the parent halo for reconstruction and integrate them in a cosmological setting (Sec. 2.4).
@@ 2.1 Simulations and coordinate system
Snapshots are saved approximately every 25 Myr over the last 7 Gyr of the simulation; the the most massive satellite merger in the last 6 Gyr has a total mass ratio of 1:45 relative to the MW (Arora et al., 2022; Garavito-Camargo et al., 2023). Sanders et al. (2020) demonstrated that reconstructed orbits in a DM-only simulation, with potential models sampled over intervals of 10 Myr and 40 Myr, yielded essentially identical results. Our sampling interval is sufficiently high to ensure high-fidelity orbit reconstructions for halo star orbits. We also show high fidelity can be achieved with a lower snapshot save rate of about 100 Myr. Additionally, the frequent snapshot intervals allow for the tracking of stellar orbits.
These zoomed-in cosmological simulations are run in an arbitrary box frame with a non-zero total momentum. Initially, we recenter the simulation onto the host galaxy frame using the iterative shrinking spheres method (Power et al., 2003) to find a center-of-mass (COM) at each time step using star particles. This halo center in comoving coordinates is defined as $\vec{x}_{\textrm{COM}}(t)\equiv\vec{r}_{\textrm{COM}}(t)/a(t)$, where $a(t)$ is the cosmological scale factor, $\vec{r}_{\textrm{COM}}(t)$ is the physical COM position.
Subsequently, we rotate all snapshots of the simulation to align the galactic disk with the XY plane at the present day–the principal axes. We define the host-centered rotation in the principal axes as the galactocentric frame. Wang et al. (2020) and Arora et al. (2022) assessed the validity of this approach, considering the constancy of the disk’s angular momentum over the past 7 Gyr of the simulations and the effectiveness of the potential modeling techniques in such systems. In m12i, the disk plane rotates by about 20 degrees over 7 Gyr up to the present day. This fixed rotation approximation also eliminates non-inertial forces associated with time-varying rotation (see Appendix. A). While mergers with massive satellites such as the LMC (Besla et al., 2007) can affect the orientation of the disk (Baptista et al., 2023), our choice of m12i, which has a quiescent merger history over the last 6 Gyr, ensures that the disk’s orientation remains stable. However, careful consideration in modeling the disk is required if the orientation changes rapidly (see Appendix A).
@@ 2.2 Potential Models
We employ the time-evolving low-order multipole potenital (TEMP) model introduced in Arora et al. (2022), fitted using a combination of BFE on the host density at each time step without assuming any symmetry conditions, constructed using AGAMA (Vasiliev, 2019b). These expansions are formulated as separable functions of 3D radial distance ($r$) or the cylindrical radius and height ($R$ and $Z$), and angular dependence, represented by orthogonal functions. To model the DM halo and hot gas ($T_{gas}\geq 10^{4.5}$ K), we use a spherical harmonic expansion in spherical coordinates to model the angular dependence ($\theta,\phi$), while the radial dependence ($r$) in the density is captured by evaluating the expansion coefficients in 25 logarithimically-spaced 1D radial grid nodes, interpolated using quintic splines (Vasiliev, 2013). The potential is written as
@@
We model the flattened stellar and cold gas ($T_{gas}\leq 10^{4.5}$ K) component using a Fourier harmonic expansion in cylindrical coordinates ($R,\phi,Z$), where the expansion coefficients are computed on a 2D meridional plane ($R,Z$) with 25 and 40 grid nodes in $R$ and $Z$ respectively. The potential is written as
@@
The methodology for fitting these models and computing the expansion coefficients ($\Phi_{\ell m}(r)$ and $\Phi_{m}(R,Z)$) is detailed in Vasiliev (2019b), and specifically applied to our simulations in Arora et al. (2022). These BFE adequately capture deformations in the disk (e.g. Wang et al., 2020; Petersen et al., 2022) and halo (e.g., Vasiliev et al., 2021; Garavito-Camargo et al., 2021) resulting from galactic evolution and satellite mergers. BFE has proven effective in reproducing orbits, even in the presence of massive satellites, in both idealized (Lilleengen et al., 2023; Vasiliev, 2023) and cosmological simulations (Lowing et al., 2011; Sanders et al., 2020; Arora et al., 2023; Donlon et al., 2024).
@@ 2.3 Force reconstruction
Note. $\mu$, $\sigma$ represent the mean and standard deviation, respectively, for each force residual distribution in different Cartesian directions (as in bottom row of Figure 1). $\mathbf{F}_{\textrm{res, tot}}$ denotes the residual on the total force magnitude.
@@ 2.3.1 Reconstructions under imposed model symmetry
Figure 2: Variation in performance with different symmetry assumptions. Force residual or the fractional variation from the true force distributions of the absolute magnitude of force for DM particles (green) and star particles (orange) within 50 kpc of the galactic center at the present day. The forces are reconstructed using fixed pole order $(\ell,m)_{\mathrm{max}}=4$ and different symmetry conditions for the halo and the disk, as denoted by the notation on the x-axis (symmetry on the halo, symmetry on the disk). The corresponding symmetry names and notation are provided in Table 2. The horizontal dotted lines represent the 25th and 75th quartiles of the distributions, while the dashed line represents the 50th quartile. “n” is a shorthand for no symmetry, “t” for triaxial symmetry, and “a” for axisymmetry.
@@
2.4 Selection of stars and orbit integration
Figure 3: Example of orbit reconstruction. Real (solid white) and reconstructed (dashed cyan) trajectories for a sample star particle extracted from the simulation and integrated in the time-evolving multipole model from Sec. 2.2 for approximately 3.8 Gyr up to the present day. The background displays the surface density of m12i in the XZ plane. The trajectories exhibit a close match, indicating a strong agreement between the real and reconstructed paths; positions at present day (i.e. the final timestep of integration) are marked with a star. Fig. 4, below, shows a sample of 5 other randomly selected orbits with varying orbital periods. All the trajectories are shown in physical coordinates.
We select a sample of about 3000 stars with tracked halo-like orbits and integrate them forward in time for approximately 4 Gyr in the TEMP model with no symmetries imposed, described in Sec 2.2. Stars in our sample are:
@@ 2.4 Selection of stars and orbit integration
We employ a leapfrog algorithm (Van Gunsteren & Berendsen, 1988) to integrate sample orbits over the last 4 Gyr of the simulation using our time-dependent potential model. The models are saved at discrete time points, so we calculate the force experienced by a test particle at any given time using linear interpolation between forces from adjacent snapshots. Each particle orbit employs a time-stepping scheme tailored to its trajectory. Initially with a time step of $\Delta t=1$ Myr, the step is scaled such that changes in velocity at each step are small relative to the velocity magnitude. This ensures resolution of pericenters for each orbit.
Fig. 3 plots the real (solid white) and reconstructed (dashed cyan) trajectory of a sample star integrated for approximately 3.8 Gyr to the present day, in the XZ plane. Both the trajectories start from the same point and the final positions at present day are marked with a star. Overall, the orbit exhibits a close match, indicating a strong agreement between the real and reconstructed paths.
@@ 3 Accuracy of Reconstructed Orbits
In this section, we quantify the accuracy of reconstructed orbits for the sample of $\sim 3000$ stars described in Sec. 2.4 with our TEMP model consisting of a spherical harmonic expansion for the DM halo and azimuthal expansion for the galactic disk (see Sec. 2.2). We quantify the spatial and temporal dependence of relative error between true and reconstructed positions for each star (Sec. 3.1). We also compare integrals of motion such as the total energy and angular momentum (Sec. 3.2), which are robust proxies for quality of our reconstructed orbits and statistically quantify “failure” modes based on a 100% error in recovering total angular momentum for an orbit.
@@ 3.1 Relative position error metric
Fig. 4 shows a selection of 5 other randomly chosen orbits, arranged by increasing orbital period. The trajectories are presented in Cartesian planes across the first three columns, with the final column illustrating the distance from the center as a function of time. Upon visual examination, the majority of orbits are reproduced with a reasonable level of accuracy. However, orbits with multiple orbital periods and an average orbital distance within 20 kpc exhibit higher errors, exceeding 10%. Additionally, deviations in other trajectories may result from interactions with substructures that are not adequately resolved with our low order expansion $(\ell,m)_{\textrm{max}}=4$. It has been suggested that subhalos on the order of a few kiloparsecs in size would require a significantly higher harmonic order (Lowing et al., 2011).

### v2

@@ Abstract
software: IPython (Perez & Granger, 2007), Matplotlib (Hunter, 2007), Numpy (Harris et al., 2020), Gizmo Analysis (Wetzel & Garrison-Kimmel, 2020a), Agama (Vasiliev, 2019b), Rockstar (Behroozi et al., 2012), Halo Analysis (Wetzel & Garrison-Kimmel, 2020b), CMasher (van der Velden, 2020).
@@ 1 Introduction
The evolution of the halo’s structure can be captured by using Basis Function Expansions (BFE) models, which can accurately describe any arbitrary density field (e.g Weinberg, 1999; Lowing et al., 2011; Vasiliev, 2013; Sanders et al., 2020; Garavito-Camargo et al., 2021; Vasiliev et al., 2021; Arora et al., 2022; Petersen et al., 2022; Lilleengen et al., 2023). Lowing et al. (2011) showed that by applying a Hernquist BFE (often referred to in the literature as a self-consistent field expansion) to every snapshot, spaced roughly 25 Myr, from a cosmological simulation of a DM halo and then interpolating the coefficients of the expansion in time allows substantial improvement in the orbital properties of a halo. Similarly, Sanders et al. (2020) showed high-fidelity orbit reconstruction in DM-only cosmological simulations of MW mass galaxies using compact BFE representation for the time-evolving halo potential, employing spherical harmonics for angular variation and bi-orthonormal basis functions or splines for radial variation. In this paper, we extend this methodology to model fully zoomed baryonic-cosmological simulations of MW mass galaxies from the Latte suite of the FIRE-2 project (Wetzel et al., 2016, 2023). Unlike Sanders et al. (2020), where only DM was considered, our simulations include stars and gas, necessitating adaptation of the BFE method with an azimuthal harmonic expansion to account for the baryonic component’s flattened shape. Additionally, we utilize a fiducial temporal cadence with snapshots spaced at approximately 25 Myr intervals, and explore the effect of larger sampling intervals (up to 500 Myr) on the quality of recovered orbits. We demonstrate that sufficiently high fidelity halo orbit reconstruction can be achieved with our fiducial cadence, even though the potential in the inner regions can change much more rapidly. We assess the efficacy of this modified potential model in reconstructing orbits at a range of radii. We fit BFE to the potential at a series of discrete snapshots. Then, we approximate the time dependence by calculating a force/acceleration on a star at a given point in time as a linear interpolation of the forces in the two time-adjacent snapshots for orbit recovery.
@@
In Sec. 2, we provide a brief overview of the simulation (§2.1), potential modeling techniques (§2.2), and constraints on recovered particle forces in the potential model (§2.3). In Sec. 3, we describe how we select orbits of halo stars from the parent halo for reconstruction and integrate them in a cosmological setting. In Sec. 4, we statistically quantify the quality of reconstructed orbits based on recovered 3D positions (§4.1), approximate integrals of motion, and orbital parameters such as pericenter and apocenter distances (§4.2). We also evaluate the dependence of orbit quality on the sampling interval for potential models (§4.4). In Sec. 5, we demonstrate an application of the model by simulating the tidal disruption of a dwarf satellite. We discuss our findings and conclusions in Sec. 6.
@@ 2 Methods
In this section, we detail our approach to reconstructing halo star orbits in a simulation of a MW-mass galaxy. We describe the galaxy simulation used (Sec. 2.1), the BFE-based potential model fits to the simulation (Sec. 2.2) and their force reconstruction for a sample of DM and star particles from the parent halo at present day (Sec. 2.3).
@@ 2.1 Simulations and coordinate system
Snapshots are saved approximately every 25 Myr over the last 7 Gyr of the simulation; the most massive satellite merger in the last 6 Gyr has a total mass ratio of 1:45 relative to the MW (Arora et al., 2022; Garavito-Camargo et al., 2024). Sanders et al. (2020) demonstrated that reconstructed orbits in a DM-only simulation, with potential models sampled over intervals of 10 Myr and 40 Myr, yielded essentially identical results. Our sampling interval is sufficiently high to ensure high-fidelity orbit reconstructions for halo star orbits. We also show high fidelity can be achieved with a lower snapshot save rate of about 100 Myr. Additionally, the frequent snapshot intervals allow for the tracking of stellar orbits.
These zoomed-in cosmological simulations are run in an arbitrary box frame on an expanding background with a non-zero total momentum. Initially, we recenter the simulation onto the host galaxy frame using the iterative shrinking spheres method (Power et al., 2003) to find a center-of-mass (COM) at each time step using star particles. This halo center in comoving coordinates is defined as $\vec{x}_{\textrm{COM}}(t)\equiv\vec{r}_{\textrm{COM}}(t)/a(t)$, where $a(t)$ is the cosmological scale factor, $\vec{r}_{\textrm{COM}}(t)$ is the physical COM position.
Subsequently, we rotate all snapshots of the simulation to align the galactic disk with the XY plane at the present day–the principal axes. We define the host-centered rotation in the principal axes as the galactocentric frame. Wang et al. (2020) and Arora et al. (2022) assessed the validity of this approach, considering the constancy of the disk’s angular momentum over the past 7 Gyr of the simulations and the effectiveness of the potential modeling techniques in such systems. In m12i, the disk plane rotates by about 20 degrees over 7 Gyr up to the present day. This fixed rotation approximation also eliminates non-inertial forces associated with time-varying rotation (see Appendix. A). While mergers with massive satellites such as the LMC (Besla et al., 2007) can affect the orientation of the disk (Baptista et al., 2023), our choice of m12i, which features a quiescent merger history over the last 6 Gyr with the most massive merger being with a satellite of mass ratio 1:45, ensures that the disk’s orientation remains stable. The models presented in this study are constructed to represent roughly the last 4 Gyr of the halo evolution, but they can be extended to any time frame as long as the disk remains oriented in the same direction. However, careful consideration in modeling the disk is required if the orientation changes rapidly, as discussed in Appendix A.
@@ 2.2 Potential Models
We employ the time-evolving low-order multipole potenital (TEMP) model introduced in Arora et al. (2022), fitted using a combination of BFE on the host density at each time step without assuming any symmetry conditions, constructed using AGAMA (Vasiliev, 2019b). We fit these expansions to each snapshot in the final 4 Gyr of halo evolution to all the particles within 600 kpc of the galactic center.These expansions are formulated as separable functions of 3D radial distance ($r$) or the cylindrical radius and height ($R$ and $Z$), and angular dependence, represented by orthogonal functions. To model the DM halo and hot gas ($T_{gas}\geq 10^{4.5}$ K), we use a spherical harmonic expansion in spherical coordinates to model the angular dependence ($\theta,\phi$), while the radial dependence ($r$) in the density is captured by evaluating the expansion coefficients in 25 logarithimically-spaced 1D radial grid nodes, interpolated using quintic splines (Vasiliev, 2013). The potential is written as
@@
We model the stellar bulge and flattened stellar and cold gas ($T_{gas}\leq 10^{4.5}$ K) component using a Fourier harmonic expansion in cylindrical coordinates ($R,\phi,Z$), where the expansion coefficients are computed on a 2D meridional plane ($R,Z$) with 25 and 40 grid nodes in $R$ and $Z$ respectively. The potential is written as
@@
The methodology for fitting these models and computing the expansion coefficients ($\Phi_{\ell m}(r)$ and $\Phi_{m}(R,Z)$) is detailed in Vasiliev (2019b), and specifically applied to our simulations in Arora et al. (2022). The BFE adequately capture deformations in the disk (e.g. Wang et al., 2020; Petersen et al., 2022) and halo (e.g., Vasiliev et al., 2021; Garavito-Camargo et al., 2021) resulting from galactic evolution and satellite mergers. BFE has proven effective in reproducing orbits, even in the presence of massive satellites, in both idealized (Lilleengen et al., 2023; Vasiliev, 2023) and cosmological simulations (Lowing et al., 2011; Sanders et al., 2020; Arora et al., 2024; Donlon et al., 2024).
@@ 2.3 Force reconstruction
Note. $\mu$, $\sigma$ represent the mean and standard deviation, respectively, for each force residual or the fractional variation from the true force distribution in different Cartesian directions (as in bottom row of Figure 1). $\mathbf{F}_{\textrm{res, tot}}$ denotes the residual on the total force magnitude.
@@ 2.3.1 Reconstructions under imposed model symmetry
Figure 2: Variation in performance with different symmetry assumptions. Force residual or the fractional variation from the true force distributions of the absolute magnitude of force for DM particles (green) and star particles (orange) within 50 kpc of the galactic center at the present day. The forces are reconstructed using fixed pole order $(\ell,m)_{\mathrm{max}}=4$ and different symmetry conditions for the halo and the disk. The corresponding symmetry names and notation are provided in Table 2. The horizontal dotted lines represent the 25th and 75th quartiles of the distributions, while the dashed line represents the 50th quartile.
@@
3 Selection of stars and orbit integration
Figure 3: Example of orbit reconstruction. Real (black solid curve) and reconstructed (blue dashed curve) trajectories for a randomly selected star particle extracted from the simulation and integrated in the time-evolving multipole model from Sec. 2.2 for approximately 3.8 Gyr up to the present day in the XY (left) and RZ (right) plane. The trajectories exhibit a close match, indicating a strong agreement between the real and reconstructed paths; positions at present day (i.e. the final timestep of integration) are marked with a star. Fig. 4, below, shows a sample of 5 other randomly selected orbits with varying orbital periods. All the trajectories are shown in physical coordinates.
To explore a representative selection of halo orbits, we select star particles in the simulation that meet the following criteria for tracked halo-like orbits:
@@ 3 Selection of stars and orbit integration
Out of approximately 10 million stars in the simulation, about 3000 fulfill these criteria. These selected stars are then integrated forward in time for approximately 4 Gyr in the TEMP model with no imposed symmetries, described in Sec 2.2. The choice of a 4 Gyr integration period is arbitrary in absolute terms, but it ensures that the orbits undergo sufficient dynamical evolution, capturing between 1 and 100 orbital periods.
@@
We employ a leapfrog algorithm (Van Gunsteren & Berendsen, 1988) to integrate sample orbits over the last 4 Gyr of the simulation using our time-dependent potential model. The integration follows eq. 1 where the terms on the right-hand side represent:
  • Gravitational force: The first term, $-\vec{\nabla}_{\vec{r}}\Phi(\vec{r},t)$, represents the gravitational force, which is calculated from the time-dependent potential model $\Phi(\vec{r},t)$. The potential models are saved at discrete time points, and at each integration step, we compute the force experienced by a particle using linear interpolation between adjacent snapshots of the potential.
  • Cosmological expansion: The second term, $-\frac{\dot{a}(t)}{a(t)}\vec{v}_{\mathrm{pec}}$, accounts for the effect of cosmological expansion, where $a(t)$ is the scale factor and $\dot{a}(t)$ is its time derivative (the Hubble parameter). We compute this term by evaluating the time-dependent scale factor $a(t)$ at each integration step, applying a velocity “kick” corresponding to the cosmological effect on peculiar velocities.
  • Motion of the center-of-mass: The third term, $-\frac{d\vec{u}_{\textrm{COM}}}{dt}$, captures the acceleration due to the motion of the galaxy’s center-of-mass (COM). To compute this term, we use spline fits to the time series of the COM velocity, $\vec{u}_{\textrm{COM}}$, for the galaxy at each snapshot. These spline fits provide a smooth representation of the COM motion, allowing us to calculate the velocity change (kick) imparted to each particle at every time step.
For each particle, the leapfrog integration carefully chooses a timestep ($\Delta t$) to resolve the pericenter of the orbit. The $\Delta t$ is computed based on the initial conditions of the orbit using the present-day potential to estimate the pericenter, ensuring sufficient resolution near critical points as:
$\Delta t=\frac{\text{tolerance}\times v_{p}}{a_{p}}$ (4)
where $v_{p}$ is the velocity at pericenter and $a_{p}$ is the acceleration at pericenter, both derived from the potential. This guarantees that the pericenter is well-resolved, preventing integration errors from accumulating in close encounters. The fixed timestep ensures that the integration errors from the leapfrog method are small compared to the force approximation errors introduced by the potential model. As a symplectic integrator, leapfrog bounds errors over time, making it well-suited for long-term orbit stability (Yoshida, 1990; Preto & Tremaine, 1999). In contrast, force errors from the potential model accumulate and dominate the overall inaccuracy.
Fig. 3 plots the real (solid black) and reconstructed (dashed blue) trajectory of a randomly selected star from our sample integrated for approximately 3.8 Gyr to the present day, in the XY (left) and RZ (right) plane. Both the trajectories start from the same point and the final positions at present day are marked with a star. Overall, the orbit exhibits a close match, indicating a strong agreement between the real and reconstructed paths.
@@ 4 Accuracy of Reconstructed Orbits
In this section, we quantify the accuracy of reconstructed orbits for the sample of $\sim 3000$ stars described in Sec. 3 with our TEMP model consisting of a spherical harmonic expansion for the DM halo and azimuthal expansion for the galactic disk (see Sec. 2.2). We quantify the spatial and temporal dependence of relative error between true and reconstructed positions for each star (Sec. 4.1). We also compare integrals of motion such as the total energy and angular momentum (Sec. 4.2), which are robust proxies for quality of our reconstructed orbits and statistically quantify “failure” modes based on a 100% error in recovering total angular momentum for an orbit.
@@ 4.1 Relative position error metric
Fig. 4 shows a selection of 5 other randomly chosen orbits, arranged by increasing orbital period. The trajectories are presented in Cartesian planes across the first three columns, with the final column illustrating the distance from the center as a function of time. Upon visual examination, the majority of orbits are reproduced with a reasonable level of accuracy. However, orbits with multiple orbital periods and an average orbital distance within 20 kpc exhibit higher errors, exceeding 10%. Additionally, deviations in other trajectories may result from interactions with substructures that are not adequately resolved with our low order expansion $(\ell,m)_{\textrm{max}}=4$. Subhalos on the order of a few kiloparsecs in size require a significantly higher harmonic order (Lowing et al., 2011).